Write an equation in slope-intercept form for the line that has the indicated slope, m, and contains the given point.
M=-1/2, (8,1)

Answer by Alan3354(69443)

Write an equation in slope-intercept form for the line that has the indicated slope, m, and contains the given point.
M=-1/2, (8,1)
====================
Use y - y1 = m*(x - x1) where (x1,y1) is the point.